# NormAd: A Benchmark for Measuring the Cultural Adaptability of Large Language Models

### v5

@@ Abstract
The integration of large language models (LLMs) into various global cultures fundamentally presents a challenge: LLMs must navigate interactions, respect social norms, and avoid transgressing cultural boundaries. However, it is still unclear if LLMs can adapt their outputs to diverse cultural norms. Our study focuses on this aspect. We introduce NormAd, a novel dataset, which includes 2.6k stories that represent social and cultural norms from 75 countries, to assess the ability of LLMs to adapt to different granular levels of socio-cultural contexts such as the country of origin, its associated cultural values, and prevalent social norms. Our study reveals that LLMs struggle with cultural reasoning across all contextual granularities, showing stronger adaptability to English-centric cultures over those from the Global South. Even with explicit social norms, the top-performing model, Mistral-7b-Instruct, achieves only 81.8% accuracy, lagging behind the 95.6% achieved by humans. Evaluation on NormAd further reveals that LLMs struggle to adapt to stories involving gift-giving across cultures. Due to inherent agreement or sycophancy biases, LLMs find it considerably easier to assess the social acceptability of stories that adhere to norms than those that deviate. Our benchmark measures the cultural adaptability (or lack thereof) of LLMs, emphasizing the potential to make these technologies more equitable and useful for global audiences. We release the dataset and code on GitHub  [^1] https://github.com/Akhila-Yerukola/NormAd.
@@ 1 Introduction
Large language models (LLMs) have become globally widespread, engaging millions of users from diverse contexts and cultures. However, studies consistently highlight cultural biases in LLM outputs, [^2] We maintain that LLMs do not inherently possess human values; however, their outputs may display knowledge and an ability to reason with certain values over others. particularly concerning the representation of various demographics (Bender et al., 2021), human values, and cultures (Masoud et al., 2023). For LLMs to be inclusive and effective across diverse and evolving cultures at scale, the model outputs must embody pluralistic values and adapt to users’ cultural nuances (Benkler et al., 2023; Rao et al., 2023). Failure to do so may lead disproportionate quality of service, cultural alienation, and a perceived lack of empathy (Wenzel and Kaufman, 2024; Lissak et al., 2024; Ryan et al., 2024).
Prior work has scrutinized language models for their knowledge of sociocultural norms. For instance, EtiCor Dwivedi et al. (2023) analyzes models’ knowledge over many societal norms across cultures. While such directive probing strategies may provide a picture of LLMs’ cultural understanding, we maintain that true multiculturalism requires models to be flexible and adjust to evolving societal and cultural norms. Molinsky (2007) highlight the benefit of cultural ‘code-switching’ among humans, adapting to different norms despite being geared to a specific set of cultural attributes. A language model should adapt to the diverse cultural settings and values it encounters. Prior works do not measure LLMs’ ability to align with such statements: As demonstrated in Figure 1, while current language models adeptly gauge the social acceptability of ‘eating with your left hand’, they struggle to align varying degrees of the same norm to user-specific realistic scenarios.
To bridge this gap, we introduce NormAd, a new benchmark designed to evaluate the cultural adaptability of LLMs. Grounded in the rich multicultural norms surrounding from Cultural Atlas (Evason et al., 2024), we’re using these social etiquette norms as a “proxy” for culture in our work. We constructed 2.6k stories that operationalize cultural norms from 75 countries to describe everyday scenarios. We filter these norms into fine-grained Rule-of-Thumb, abstracted Value paradigms, and Country name, as shown in Figure 1. Each story in our benchmark includes QA pairs for assessing social acceptability under these different cultural contexts. These questions, coupled with varying cultural contextualization degrees, enable us to evaluate models’ adherence versatility.
@@
Through comprehensive experiments with open and closed source models on NormAd, we reveal several important findings (§5): (1) Existing models struggle to answer social acceptability questions across various contextualization levels in stories, especially concerning values and country contexts. The best performing models, GPT-3.5-turbo and Mistral-Instruct, achieve  60% accuracy for Value and  55% for Country contexts. Even with all necessary information (Rule-of-Thumb), the best performing models, GPT-4footnote 5 at 87.6% and Mistral-Instruct at 81.8% perform decently but lag behind human performance (95.6%), leaving room for improvement, (2) Models struggle significantly in answering social acceptability questions involving stories that violate or are irrelevant to certain cultural social norms, suggesting the presence of agreement or sycophancy biases in models, (3) While an increase in the number of model parameters or adopting an better preference tuning optimization method helps improve overall performance, these improvements show greater performance gains in stories revolving around English-speaking and European countries (like USA) than in stories around African-Islamic cultures (like Saudi Arabia).
Overall, our work shows that current LLMs struggle with adhering to cultural norms. This highlights the need for improved contextualization capabilities in LLMs, particularly in terms of adherence and cultural adaptability. The global deployment of LLMs emphasizes the importance of ensuring the effectiveness and ethical application of language technologies in diverse cultural contexts.
Figure 1: We contrast our work from previous work in that we test for the ability of a language model to change/adapt its responses when contextualized with cultural information.
@@ 2 Related work
In §2.1, we discuss psychological studies on human ethics, values, and culture. We subsequently address existing work on LLMs evaluation strategies (§2.2) based on these theories for cultural alignment and social reasoning.
2.1 Ethics and culture - A primer
The fields of human ethics and psychology offer various theories and frameworks for understanding human perceptions within societal contexts. Kohlberg’s theory of morality Levine et al. (2009) measures the stages of moral development in humans. From a cultural standpoint, Hofstede (1980) provides a 4-dimensional cultural theory across power-distance inequalities, masculinity, uncertainty tolerance, and individualism, which is extended by Schwartz (2012) by considering ten universal human values that are shared between these cultural settings. Work has also studied culture-specific diversities; for instance, the World Values Survey (WVS) (WVS, 1981) is a long-running yearly questionnaire measuring attitudes towards societal aspects around the world, plotted on the Inglehart-Welzel Cultural map Inglehart and Welzel (2023). Each of these countries are binned into 8 cultural clusters based on their society’s historical heritage and cultural values: English Speaking, Protestant Europe, Catholic Europe, Orthodox Europe, Confucian, West and South Asia, Latin America, African and Islamic. The norms themselves are documented by the cultural atlas (Evason et al., 2024). Drawing parallels with Kohlberg’s theory and the WVS, we wish to evaluate the reasoning capabilities in LLMs across cultural contexts.
2.2 Related work: Exploring ethics and culture in Language Models
Current approaches in evaluating cultural biases have utilized the ethical frameworks and surveys derived from the aforementioned psychological theories. For instance, researchers have applied frameworks such as Hofstede’s cultural dimensions (Masoud et al., 2023), Kohlberg’s theory of morality (Tanmay et al., 2023), and Schwartz’s theory of basic values (Yao et al., 2023) to gauge the moral sensitivity of LLMs. Additionally, many culture-related work use several proxies to determine cultural acceptability (Adilazuarda et al., 2024); such as using ethical principles Hendrycks et al. (2023) or human values through World-Value Survey (WVS) data (Johnson et al., 2022; Atari et al., 2023; Masoud et al., 2023; AlKhamissi et al., 2024; Ramezani and Xu, 2023). Several other works prefer real-world social norms, probing a model’s world knowledge (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024; Dwivedi et al., 2023), with some attempting to instill social norms directly into language models (Dwivedi et al., 2023) or finetune a multilingual LLM over a WVS sample and show performance gains over downstream tasks such as hate speech detection (Li et al., 2024). However, finetuning or directive probing methodologies face limitations, as much of the focus tends to be on knowledge acquisition rather than infusing reasoning capabilities. To the best of our knowledge, no work has attempted to measure cultural adaptability or flexibility of language models – we distinctively measure the adaptability of LLMs through their degree of applicability of cultural norms over social situations.
3 NormAd construction
In order to investigate the adaptability of LLMs to multifaceted social etiquette-relate norms across different cultures, we introduce NormAd. Leveraging cultural information from 75 countries, we use an automated human-in-the-loop generation process to construct narrative stories. These stories depict everyday interactions between characters, involving fine-grained Rules of Thumb (RoT), abstracted Value paradigms and Country-specific social etiquette. We use question-answer pairs to measure social acceptability of constructed stories under these different degrees of adaptability. In this section, we describe our three-step data construction pipeline: (1) narrative generation, (2) filtration, and (3) validation.
Figure 2: The story generation process. We source stories from the cultural atlas through a generative process, followed by automated and manual validation.
3.1 Narrative Generation
3.1.1 Data Sourcing
We collect data about various cultural norms across 75 different countries from the Cultural Atlas (Evason et al., 2024) [^3] https://culturalatlas.sbs.com.au. The Cultural Atlas, launched by multiple Australian organizations, aims to “inform and educate the (Australian) public in cross-cultural attitudes, practices, norms, behaviors, and communications". We select this as our cultural data source, as it includes global community interviews and rigorous validation by community experts and academic researchers [^4] The multiple validation stages for the norms have been detailed here. Using the taxonomy from Adilazuarda et al. (2024), we focus on measuring culture through a semantic proxy with social norms, focusing on the ‘Etiquette’ category specifically as it covers socially acceptable and unacceptable norms in scenarios such as dining or visiting homes. We divide this into four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gift-Giving’. Each subcategory comprises 5-10 socially endorsed or discouraged norms specific to each country.
3.1.2 Synthetic story generation
Using the above data source as seeds, we construct “cultural stories” representing the cultural norms of each country. We adopt the use of LLM-grounded generation to help depict realistic social interactions in the world (Gordon and Van Durme, 2013; Tan et al., 2024). Each story involves an everyday situation, with a question regarding its social acceptability. Social acceptability depends on the cultural context of the situation, such as the location it takes place in and its associated social norms. Hence, we define three forms of cultural contexts needed to answer social acceptability questions:
Rule-of-Thumb
(RoT) represents all the nuanced information about a social norm necessary to answer questions about social acceptability of a character in a story. Note that it is stripped of all geographical information. Ideally, humans and models should be able to easily evaluate social acceptability questions of character actions within a story based on this alone, as this simplifies to an entailment task.
@@ 3.1.2 Synthetic story generation
indicates the country in which the social situation in the story occurs. This allows us to focus on the LLM’s ability to apply its internal or external retrieval-based knowledge of a social norm associated with the country.
Value
We also consider an abstraction of the rule-of-thumb, representing broader human principles surrounding specific social norms. This is a middle-ground between knowledge and adaptability measurement: we wish to test models’ ethical reasoning by their ability to apply broader/shared cultural principles to specific situations, potentially leveraging some of their intrinsic knowledge.
By employing this methodology, we can delineate a hierarchy/degree of contexts: [1] Country, which hinges on an LLM’s intrinsic knowledge of norms; [2] Value, which introduces an abstraction surrounding a social norm; and [3] RoT, which presents the norm itself. We create synthetic stories and their corresponding contexts using gpt-4-turbo. We provide gpt-4-turbo with the Country of origin and its cultural background from the Cultural Atlas as context, and we carefully prompt it to generate a narrative story, along with the RoT and Value encompassing it. The generated stories vary in acceptability, violation, and irrelevance to these norms (§3.1.3). By ensuring that no cultural knowledge can be directly inferred from the story, through the exclusion of geographical indications such as country names, we force the models being tested to align with the provided contextualization when answering questions. We provide few-shot examples of fictional stories and contexts that follow the aforementioned properties to gpt-4-turbo, which are described in Appendix B.5. Example cultural stories and contexts from NormAd are provided in Table 1.
3.1.3 Story Answers
Our generated stories follow three groups based on adherence to social norms:
Adhering to Social Norm (Yes)
@@ 3.1.3 Story Answers
Stories generated in this category align characters’ actions with known social norms or etiquette of their cultural backgrounds. For example, if the cultural norm dictates using the right hand for certain actions, the generated story would include characters performing those actions with their right hand.
Violating a Social Norm (No)
@@
Here, stories depict deviations or violations of the established social norms of the cultural context. This is achieved by scripting scenarios where characters engage in actions that are considered culturally inappropriate based on the known norms.
Neutral Situation (Neutral)
These stories neither adhere to nor violate a given social norm. They are contextualized from irrelevent backgrounds outside the native country’s Inglehart-Welzel cluster. We ensure their neutrality in social acceptability through a validation process detailed in §3.2.
Table 1 provides example stories for each group and context from NormAd. Each story’s assigned group is treated as its ground-truth ‘label’. We generate three stories for every (norm, story label) combination present in the etiquette category of Cultural Atlas. Upon providing a story to an LLMs, we elicit an answer from it across these three labels. The prompt is specified in Appendix A.3.
@@
Story At a dinner gathering, Alice offered Mark a plate of food using her right hand. Is what Alice did socially acceptable? At a dinner party at his friend Tom’s house, Michael eagerly finished his plate and made sure it was completely empty, wanting to show how much he enjoyed the meal. Is what Michael did acceptable? Alex was having lunch at a friend’s house. When another friend arrived unexpectedly, Alex immediately offered some of their meal to the newcomer. Is what Alex did socially acceptable?
@@
Table 1: Sample stories with contextualizations and labels from NormAd
3.2 Story Filtration
To ensure validity and quality of the generated stories and their corresponding RoT and Value, we conduct several rounds of automated tests:
Check 1: Validate relevance of RoT to stories
We manually inspect a sample of 200 generated stories and their corresponding RoT to check if the RoT is relevant to its corresponding story. For the ‘Neutral’ label, we observe some inconsistencies and use gpt-4 to filter and regenerate the inconsistent stories. This process is repeated three times, and we discard stories that do not conform to the ’Neutral’ gold-label.
Check 2: Entailment of RoT to cultural background
We use gpt-4 to verify the relevance/entailment of the Rule-of-Thumb to its corresponding country’s background (excluding ’Neutral’ labeled stories, where we ensure relevance to different country’s background). We identified, verified and discarded 173 stories without an aligned Rule-of-Thumb.
Check 3: Ensure Value is an abstraction of RoT using entailment
We use gpt-4 to verify if Value entails a RoT. We identified a very small number of stories ($\sim 20$) that were misaligned, and manually correct the RoT and Values.

### v6

# NormAd: A Framework for Measuring the Cultural Adaptability of Large Language Models

@@ Abstract
To be effectively and safely deployed to global user populations, large language models (LLMs) must adapt outputs to user values and culture, not just know about them. We introduce NormAd, an evaluation framework to assess LLMs’ cultural adaptability, specifically measuring their ability to judge social acceptability across different levels of cultural norm specificity, from abstract values to explicit social norms. As an instantiation of our framework, we create NormAd-Eti, a benchmark of 2.6k situational descriptions representing social-etiquette related cultural norms from 75 countries. Through comprehensive experiments on NormAd-Eti, we find that LLMs struggle to accurately judge social acceptability across these varying degrees of cultural contexts and show stronger adaptability to English-centric cultures over those from the Global South. Even in the simplest setting where the relevant social norms are provided, our best models’ performance ($<$82%) lags behind humans ($>$95%). In settings with abstract values and country information, model performance drops substantially ($<$60%), while human accuracy remains high ($>$90%). Furthermore, we find that models are better at recognizing socially acceptable versus unacceptable situations. Our findings showcase the current pitfalls in socio-cultural reasoning of LLMs which hinder their adaptability for global audiences.
NormAd: A Framework for Measuring the
Cultural Adaptability of Large Language Models
[…]
@@ 1 Introduction
Large language models (LLMs) have become globally widespread, engaging millions of users from diverse contexts and cultures Kasneci et al. (2023); Yuan et al. (2022). However, studies consistently highlight cultural biases in LLM outputs, [^1] We maintain that LLMs do not inherently possess human values; however, their outputs may display knowledge and an ability to reason with certain values over others. particularly concerning the representation of various demographics (Bender et al., 2021), human values, and cultures (Masoud et al., 2023). To be inclusive and effective across evolving cultures, LLM outputs must embody pluralistic values and adapt to users’ cultural nuances (Benkler et al., 2023; Rao et al., 2023); otherwise, there is a risk of providing disproportionate quality of service and fostering cultural alienation (Wenzel and Kaufman, 2024; Lissak et al., 2024; Ryan et al., 2024).
Previous work has largely focused on assessing knowledge and biases by probing LLMs with curated socio-cultural knowledge databases Nguyen et al. (2023); Dwivedi et al. (2023); Fung et al. (2024); Shi et al. (2024), often using direct questions about cultural norms, such as, ‘Is it okay to eat with your left hand in India?”. While these methods provide insights into what models know about different cultures, they do not fully evaluate their overall multi-cultural competence Deardorff (2009); Hovy and Yang (2021). We argue that true cultural competence requires models to not just possess cultural knowledge, but also to apply it flexibly to user-specific scenarios. Molinsky (2007) highlights the benefit of cultural ‘code-switching’, which allows humans to adapt to different norms despite being attuned to their own cultural attributes. Similarly, LLMs should be culturally adaptable, i.e., able to adjust their responses based on the user’s cultural context. By utilizing cultural values provided by or inferred from the user, rather than rigidly adhering to internal biases, these models can ensure effective communication across diverse scenarios.
Figure 1: We introduce NormAd, a framework for testing a language model’s ability to adapt its responses when contextualized with varying levels of cultural information specificity, in contrast to prior methods that directly probe models for their knowledge. We show that LMs struggle to pick up cultural cues when provided with varying levels of context, unlike humans, who can generally recognize such cues.
To address the gap in evaluating the cultural adaptability in LLMs, we introduce the NormAd evaluation framework (§3). Using social norms as a proxy for culture Adilazuarda et al. (2024), NormAd evaluates how models reason about the acceptability of social situations described in free-text, under varying levels of socio-cultural context. As shown in Figure 1, each situational description is evaluated with varying degrees of cultural norm specificity: specific Country names, an abstract high-level Value along with Country names, and fine-grained RoT. This hierarchical approach evaluates LLMs’ ability to understand and apply cultural norms, while testing their performance across varying levels of cultural context that might be provided in real-world scenarios.
As an instantiation of our framework, we develop NormAd-Eti4), a benchmark for measuring cultural adaptability specifically focused on social etiquette norms in English. These multicultural norms are sourced from Cultural Atlas (Evason et al., 2024), an educational resource built on extensive global community interviews and rigorous validation. NormAd-Eti contains 2.6k social situations from 75 countries, each with a question-answer pair to evaluate LLMs’ ability to judge the social acceptability of specific actions across various cultures and levels of cultural norm specificity.
@@
Through comprehensive experiments with open and closed source models on NormAd-Eti6), we find that: (1) Current models struggle with social acceptability questions across all levels of cultural norm specificity and contextualization, particularly in Value and Country contexts. (2) Models struggle significantly in answering questions involving situational descriptions that violate or are irrelevant to certain cultural social norms, suggesting potential agreement or sycophancy biases, (3) While increasing the number of model parameters or using better preference tuning optimization methods improves performance, these gains are more pronounced in social situations revolving around English-speaking and European norms (e.g., USA) than in those related to African-Islamic cultures (e.g., Saudi Arabia).
Through NormAd, we demonstrate LLMs’ struggle to judge social acceptability across varying cultural contexts, highlighting the critical need for better cultural contextualization capabilities. We discuss the importance, complexity, and limitations of evaluating cultural knowledge and adaptability (§8), promoting approaches, such as ours, that allow for user-provided cultural context.
@@ 2 Related work
2.1 Culture in LLMs
Recently, several studies have examined the socio-cultural reasoning of LLMs, evaluating their ability to serve diverse users and values. Some studies have used psychological and cultural surveys WVS (1981); Hofstede (1980) to prompt models with human values Johnson et al. (2022); Atari et al. (2023); Masoud et al. (2023); Ramezani and Xu (2023), guaging how well these models reflect diverse cultural values. Other studies have focused on probing LLMs for their cultural knowledge of social norms (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024). While Dwivedi et al. (2023) explored etiquette-related norms through direct probing for knowledge, our approach instead measures adaptability. Studies have also investigated LLMs’ knowledge of cultural artifacts such as food, art forms, clothing, and geographical markers Seth et al. (2024); Li et al. (2024); Koto et al. (2024). These evaluations have primarily focused on measuring the knowledge component of cultures in LLMs, rather than applying and adapting their knowledge to user-specific scenarios. Efforts to improve adaptability have mostly focused on enabling LLMs to adopt synthetic personas from different regions AlKhamissi et al. (2024); Durmus et al. (2023); Kwok et al. (2024).
Overall, these studies have helped reveal gaps in cultural knowledge, especially regarding non-western cultures, and have complemented known stereotypes and demographic biases in LLMs Bhatt et al. (2022); Zhou et al. (2022); Jha et al. (2023). Some efforts have aimed to address these issues by fine-tuning LLMs to instill social norms (Dwivedi et al., 2023) or improve performance on culture-specific tasks, such as hate speech detection (Li et al., 2024). Interestingly, several works have shown that probing LLMs in languages associated with certain cultures, counterintuitively, does not perform better than probing them monolingually in English Shen et al. (2024); Durmus et al. (2023).
2.2 On Value Pluralism and Personalization of LLMs
Cultural studies in LLMs inherently involve dealing with conflicting values, a term known as ‘value pluralism’. Several works have studied this broader problem through either benchmark datasets Ren et al. (2024); Sorensen et al. (2024a); Pistilli et al. (2024), finetuning models to respond pluralistically and prosocially Kim et al. (2022); Forbes et al. (2020) or by proposing modular frameworks around value pluralism Benkler et al. (2023); Feng et al. (2024). Our work is pluralistic in that it prompts LLMs with situations that can have potentially conflicting social acceptabilties depending on context.
Figure 2: Our NormAd-Eti construction pipeline consists of 4 parts: a) Generation: We source social etiquette-related social norms from Cultural Atlas and systematically transform them into grounded social situation description, RoT, and Value b) Filtration: We perform three rounds of automatic filtering and sanity checks to eliminate inconsistencies c) Validation: We conduct extensive human validation of the constructed dataset d) Human Performance: We conduct a small-scale assessment of human performance.
3 NormAd Evaluation Framework
We introduce a multi-level evaluation framework to measure the cultural adaptability of LLMs, contrasting existing work that primarily measures knowledge2.1). To be inclusive of diverse populations with varying values Sorensen et al. (2024b), we argue that a truly adaptable LLM should perform well across diverse cultural contexts.
Our framework centers on free-text descriptions of social situations with multiple characters, intentionally devoid of explicit cultural or geographical markers. As shown in Figure 1, each scenario includes a social acceptability question about a character’s actions. Recognizing real-world scenarios’ varying cultural information, we evaluate LLMs’ adaptability across 3 levels of cultural specificity:
Rule-of-Thumb (RoT)
Detailed information necessary to answer social acceptability questions about character actions, simplifying the task to an entailment problem for both humans and models. For instance, Figure 1 describes a situation where Sarah is eating with their left hand and the RoT is to “avoid eating with your left hand”.
@@ 3 NormAd Evaluation Framework
The country where the social situation occurs. Truly culturally adaptable LLMs should perform this task by combining knowledge of country-specific cultural norms (acquired during training or through external retrieval) with country-level contextualization. In the above example, given only that the situation takes place in “India”, the LLM should infer that eating with the left hand is generally considered disrespectful in India. We expect LLMs, unlike humans, [^2] Most humans lack complete knowledge of all cultures, as even members of a specific culture might not be familiar with every nuance and value within their own cultural context. to perform this task well across diverse cultures.
Value +Country
An easier version of the Country setting, where both an abstract high-level value derived from the RoT and the country are provided. Similar to Country setting, LLMs should infer the social norm for that Country and Value. For instance, given “hygiene in dining” and “India”, an LLM should infer the norm of not eating with the left hand based on Indian dining customs related to hand usage.
4 NormAd-Eti Construction
We demonstrate the utility of our framework by constructing NormAd-Eti to explore LLMs’ adaptability to social etiquette norms across different cultures. Grounded in the rigorously validated Cultural Atlas resource, we generate situational descriptions in English across 75 countries. In this section, we describe our data construction pipeline (see Figure 2): (1) Social situation description, (2) Automatic Filtration, (3) Human Validation, and (4) Verification of Human Performance.
4.1 Social Situation Description
Norm Sourcing
We select social-etiquette norms across 75 different countries from the ‘Etiquette’ category of Cultural Atlas (Evason et al., 2024). [^3] https://culturalatlas.sbs.com.au The Cultural Atlas, launched by multiple Australian organizations, aims to “inform and educate the (Australian) public in cross-cultural attitudes, practices, norms, behaviors, and communications". We select this as our data source, as it includes global community interviews (with translators) and rigorous validation by community experts, religious leaders, and academic researchers. [^4] We acknowledge that no singular data source will be a complete and accurate representation of the broad concept of culture. We choose this source as a proxy, primarily due the multiple validation stages for the norms, which have been detailed here.
The Etiquette category from the Cultural Atlas covers acceptable and unacceptable behaviors across various subcategories, such as dining, home visits, and giving compliments, with each subcategory containing approximately 5-10 culturally specific norms per country. These subcategories may vary or be missing in different countries. Ultimately, we regroup them into four main categories: Basic Etiquette, Eating, Visiting, and Gift-Giving.
Social Situation Labels
We construct social situation descriptions with three types of labels:
  1. 1.
@@ 4.1 Social Situation Description
    Adhering to Social Norm (Yes) Here, situation descriptions align characters’ actions with known cultural norms. For example, if a norm dictates using the right hand for certain actions, the situation would depict characters doing so.
  2. 2.
@@
    Violating a Social Norm (No) Here, situation descriptions depict deviations or violations of the established cultural norms, portraying characters engaging in culturally inappropriate actions.
  3. 3.
    Neutral Situation (Neutral) These descriptions neither adhere to nor violate a given social norm.
Transforming Norms into Social Situation Descriptions
Grounded in etiquette-related norms, Drawing inspiration from Kim et al. (2023), we systematically transform etiquette-related norms into grounded social situation descriptions. For each country and subcategory present in Cultural Atlas, we generate nine situations: three per label. We prompt gpt-4-turbo with cultural norms for each subcategory and the desired label, instructing it to generate a situational description based on the norm, along with a corresponding RoT and Value. Few-shot examples, varying by target label, are provided as well. For the ‘Yes’ and ‘No’ labels, we use cultural norms from the source country. For the ‘Neutral’ case, we select cultural norms from a different cluster on the Inglehart-Welzel cultural map Inglehart and Welzel (2012) [^5] The Inglehart-Welzel Cultural map is a plot of WVS data over traditionalist and survivalist axes. Each of these countries is binned into one of 8 clusters based on their society’s historical heritage and cultural values: English Speaking, Protestant Europe, Catholic Europe, Orthodox Europe, Confucian, West and South Asia, Latin America, African and Islamic. , outside the source country’s cluster. Importantly, via prompt instructions, we ensure that the generated descriptions does not contain any geographical indicators, such as country names. We provide the generation prompt in Appendix A.1, and a few-shot example in Appendix D.5.
This approach enables us to generate situations across diverse cultural contexts and levels of norm adherence. By excluding direct geographical references, models must rely solely on provided context, enabling a more rigorous evaluation of their understanding of cultural norms and social reasoning. See Table 1 for examples.
@@
Situation At a dinner gathering, Alice offered Mark a plate of food using her right hand. Is what Alice did socially acceptable? At a dinner party at his friend Tom’s house, Michael eagerly finished his plate and made sure it was completely empty, wanting to show how much he enjoyed the meal. Is what Michael did acceptable? Alex was having lunch at a friend’s house. When another friend arrived unexpectedly, Alex immediately offered some of their meal to the newcomer. Is what Alex did socially acceptable?
@@
Table 1: Sample situations with contextualizations and labels from NormAd-Eti
4.2 Automatic Filtration
We conduct three rounds of filtration and re-generation. We use gpt-4 to verify the relevance via entailment of the RoT with respect to situational descriptions after each round. Situational descriptions inconsistent with the gold label are regenerated in each round. The prompt is present in Appendix A.2. After three rounds, we re-assign the extra Cultural Atlas subcategories (e.g., ‘giving compliments’) into one of four designated subcategories mentioned in §4.1, resulting in 2,726 situational descriptions across 75 countries.
To further ensure the quality of the generated data, we conduct two additional automated checks, the prompts of which are in Appendix A.4:
Check 1: Entailment of RoT to Cultural Atlas’s norms
For data points with ‘Yes’ and ‘No’ gold labels, we use gpt-4 to verify if the generated RoT is derived from and relevant to the given country’s norms in Cultural Atlas. We measure this via entailment, i.e., asking gpt-4 to classify whether the country-specific norms entail the RoT. For ‘Neutral’ labels, we check if the generated RoT is irrelevant. Through this process, we identified, manually verified, and discarded 73 data points without an aligned RoT.
Check 2: Ensure Value is a high-level abstraction of RoT
We use gpt-4 to verify if Value is a high level abstraction of the corresponding RoT. Through this process, we identified and discarded 20 data points that were misaligned.
Statistics